OpenStudy (anonymous):

Find the value of k such that -2/3 is a zero of the function f(x) = {4x + k}/{7}

OpenStudy (anonymous):

@ganeshie8 Can you help me?

OpenStudy (camper4834):

8/3

OpenStudy (anonymous):

How did you get that @camper4834

OpenStudy (camper4834):

well... if you plug in -2/3 into x you get (4)(-2/3)+(k)/(7)=0

OpenStudy (camper4834):

((4)(-2/3)+(k))/(7)=0

OpenStudy (camper4834):

so -8/3 + k = 0 solve for k

OpenStudy (anonymous):

Ok thanks!

OpenStudy (camper4834):

don't forget to hit CLOSED QUESTION